OpenStudy (anonymous):

What is the simplified product of (√3)(√6)=

OpenStudy (anonymous):

\[\sqrt{a}\sqrt{b} = \sqrt{ab}\]

OpenStudy (anonymous):

9√2 √18 3√2 3

OpenStudy (anonymous):

\[3\sqrt{2}\] You got it.

OpenStudy (anonymous):

I do not know which answer it is

OpenStudy (anonymous):

Oh, well look here. You see my first post?

OpenStudy (anonymous):

3√2

OpenStudy (anonymous):

Here: \[\sqrt{3}\sqrt{6} = \sqrt{3 * 6}\]

OpenStudy (anonymous):

Now what do you have?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

√3+ √12

OpenStudy (anonymous):

that is my simplified answer

OpenStudy (anonymous):

4.24175

OpenStudy (anonymous):

No, please stop doing this. That is not what we are looking for.

OpenStudy (anonymous):

Follow me step by step. What is: \[\sqrt{6 \times 3}\]

OpenStudy (anonymous):

This is my last assignment for this term.................I do appreciate the help honestly I do.

OpenStudy (anonymous):

what mr.qpha?

OpenStudy (anonymous):

\[\sqrt{6 \times 3} = \sqrt{18} = \sqrt{ 3 * 3 * 2} = 3 \sqrt{2}\]

OpenStudy (anonymous):

2.44 =√6

OpenStudy (anonymous):

\[\sqrt{18}\]

OpenStudy (anonymous):

I feel so stupid

OpenStudy (anonymous):

Don't ever feel stupid. It just takes seeing it over and over to realize what's happening. It's building blocks. I was in the same place, I promise you that ;)

OpenStudy (anonymous):

well I am almost 50yrs old and never learned algebra in high school

OpenStudy (anonymous):

so did I get it right?

OpenStudy (anonymous):

I have a 4.0gpa but now this term because of Algebra

OpenStudy (anonymous):

Yes, but sqrt(18) simplifies though. What I do is called a "prime factorization".

OpenStudy (anonymous):

not this term that is

OpenStudy (anonymous):

I cannot thank you enough!

OpenStudy (anonymous):

I have more.................problems to solve

OpenStudy (anonymous):